# Automatically Detecting Checked-In Secrets in Android Apps: How Far Are We?

### v1

@@ 5.2.1. Complimentary among the tools
In contrast, machine-learning-based tools, in general, can adapt to a wider range of cases since they do not require any pre-defined patterns. However, our results suggest that simply providing more training sets cannot yield a significant improvement. The effectiveness of the context model is significantly degraded by the presence of obfuscation. Relying on code snippets as input for the model is proven to be ineffective in our evaluation. This finding suggests the need for a new feature extraction approach that is obfuscation-resilient.
5.2.2. String Group Analysis
@@
After analyzing the methods that contain a checked-in secret, we propose to use string groups as features to detect checked-in secrets. A string group refers to all the strings within one single method. In our manual inspection, we observed that string groups that contain checked-in secrets shared some similar patterns. For instance, the names of cloud providers like ”Google” or ”Firebase,” along with terms related to secrets such as ”user,” ”token,” or ”key,” frequently appear in these groups. This suggests that the patterns in string groups could serve as potential indicators of checked-in secrets. Moreover, since string literals are typically not modified by the obfuscation of apps in Google Play Store (Dong et al., 2018).
@@
To verify the feasibility of this approach, we conducted a preliminary study with the hypothesis that string groups containing checked-in secrets exhibit similarities to one another and are distinct from those that do not contain any secrets. To test this hypothesis, we designed a statistical analysis process. First, we collected two collections of string groups: Collection Secret included 602 string groups containing Google API keys and secrets obtained by the Three-Layer Filter. We excluded the string groups that only contain checked-in secret itself. To prevent potential bias in terms of quantity, the second set, Collection NoSecret, included 602 randomly selected string groups that did not contain any checked-in secrets. Methods with only a single string were excluded to ensure meaningful comparison. We then calculated the similarity between string groups using cosine similarity (Han et al., 2012), which is better for this task as it works with non-English text, unlike Levenshtein distance (Navarro, 2001) or Word Movers’ Distance (Saxena, 2019). Cosine similarity measures the cosine of the angle between two character vectors, generating a value between 0 and 1, where 1 indicates identical word vectors and 0 indicates no similarity. The similarity score is calculated using the equation below (Han et al., 2012).
(1) $\text{sim}(x,y)=\frac{x\cdot y}{\|x\|\cdot\|y\|}$
@@
For each string group in Collection Secret, we first transform the string groups into the character vector, in which the number in the vector represents the frequency of a character. Then, we computed the average cosine similarity both within Collection Secret (Comparison-SvS) and between Collection Secret and Collection NoSecret (Comparison-SvN). We need to verify that the two comparisons are statically different to test our hypothesis. We applied both F-test (fte, 2024) and Z-test (zte, 2024). These tests assessed whether the means and variances of the cosine similarity scores in Comparison-SvS and Comparison-SvN were significantly different. The statistical significance was determined through p-values, with smaller p-values indicating a greater likelihood that the differences in means or variances were not due to chance. Additionally, given the various forms that strings can take in Android apps, we modified Collection Secret and Collection NoSecret into three distinct versions: case-sensitive, case-insensitive, and English Word Extraction. The first two variations considered whether string case impacted the similarity calculations, while the English Word Extraction version isolated all English words using a dictionary (dolph, [n. d.]) for further analysis.
@@
The result of the two tests is shown in Table 7. The calculated p values of the two tests for the three versions are all very close to 0 and much smaller than the common significance level alpha (0.01) (alp, 2015), indicating that differences of Comparison S-S and Comparison S-N are statistically significant. We observed that the difference in the case-sensitive comparison is larger than in other variations. This is likely because ordinary strings tend to use more capital letters. Furthermore, the mean similarity score in Comparison S-S is also higher than in Comparison S-N.
@@
Such test results demonstrate the potential to leverage string groups to detect checked-in secrets in obfuscated Android apps. We plan to further explore this direction to propose a new machine-learning based checked-in secret detection technique based on string groups.
Table 7. Result of F-test and Z-test
F-Test(p value) Z-Test(p value) Similarity Score Mean (Comparison s-s: Comparison s-n)
@@
English Word Extraction $7.6\times 10^{-12}$ 0 0.65 : 0.55
6. Discussion
@@
Severity of Checked-in Secret Issue in Android apps. Most of the existing studies focused on the checked-in secrets problem on code-sharing platforms, neglecting the realm of Android applications distributed via platforms like the Google Play Store. Consequently, there exists a gap in systematic analyses of checked-in secrets issues within such projects. Our study revealed that excluding the potential false positive, 2142 unique valid checked-in secrets were found among 2,115 Android apps, which accounts for nearly half of our entire dataset. This result underscores the prevalence of the checked-in secret issues in Android apps.
Negligence of Checked-in Secret Issue. We reached out to the developers of Android apps affected by the checked-in secret issue. Some developers acknowledged the problem and committed to addressing it, while others acknowledged the issue but asserted that the checked-in secret was safe to expose. While it is true that some checked-in secrets, such as the Firebase API key, are considered safe to expose in code (though not recommended) (fir, 2024). Or Twitter Client Secret, which would only become harmful if the associated Twitter Client ID is also leaked, these keys can still pose a security threat when other related information is also exposed. For instance, if project IDs and authentication tokens associated with these checked-in secrets are leaked, they can compromise security. In our experiment, we successfully extracted additional information associated with the Firebase API key that could be harmful when combined, demonstrating the potential risk even for seemingly harmless checked-in secrets.
@@ 8. Related work
Checked-in Secret Detection for GitHub. Leakage of checked-in secret issue in the code-sharing platform is a long-studied subject. Besides the tools we selected to run the experiment, there exist other papers that also targeted the checked-in secret issue. Saha et al.  (Saha et al., 2020) introduce a machine learning-based detection tool; they designed 20 features for a string to reduce the false positive, and Lounici et al. (Lounici et al., 2021) examined the performance of various machine learning model when performing checked-in secret detection tasks. There are also some community checked-in secret detection tools. As an example, TruffleHog (tru, 2024), which was originally designed for GitHub, has also been adapted to the Android apps on the Google Play store. However, the Three-Layer Filter, which utilizes the same methodology, outperforms TruffleHog. Some researchers have also proposed tools that can detect human-generated passwords. Fadi et al.  (Sibai et al., 2008) utilized neural networks to predict whether a string is a human-generated password. While Lykousas et al. (Lykousas and Patsakis, 2023) examined the developers’ traits when choosing passwords. Those studies all aimed to develop an effective checked-in secret detection tool for open-source projects but did not examine the feasibility of their approaches on Android apps.
@@
Empircal Study on Checked-in Secrets Issue. There have also been several researches that conducted empirical studies on checked-in secret topics. Basak et al. (Basak et al., 2023a) evaluated the performance of all the available checked-in secret detection tools on a benchmark dataset (Basak et al., 2023b) that contains 818 Github repositories. They also discussed developers’ challenges about checked-in secret by analyzing 779 questions on Stack Exchange (Basak et al., 2023c). Whereas Rahman et al. (Rahman et al., 2022) aimed to investigate the phenomenon of developers disregarding warnings issued by checked-in secret detection tools, often due to the tools’ high false positive rates. Furthermore, Krause et al. (Krause et al., 2022) analyzed the reason why developers would commit checked-in secrets by interviewing other developers. Naiakshina et al.  (Naiakshina et al., 2017) studied how developers deal with checked-in secret storage through an experiment involving 20 developers. However, these empirical analyses have primarily focused on open-source projects, which may not accurately represent the situation in Android apps. We aim to address this limitation in our research.
@@
Security Concerns in Mobile Apps. The popularity of mobile apps has also drawn researchers’ attention. Many studies were conducted to analyze the security issues in mobile apps. PAMDroid (Zhang et al., 2020) was developed to investigate the privacy leakage issue caused by the misconfiguration of analytic service. DroidContext (Wang et al., 2016) targeted the detection of privacy leakage and is capable of differentiating malicious privacy identification and benign privacy disclosure identification. Bunyakiati et al. (Bunyakiati and Sammapun, 2019) examined current secret management and handling in the mobile apps development cycle. Whereas Rahat et al. developed OAUTHLINT (Rahat et al., 2019) to detect the security mistakes of the Android apps that use OAuth API. our work further investigated the security issue caused by checked-in secret.
9. Conclusion
@@
In this paper, we conducted a literature review to categorize the existing checked-in secret detection technique, selected three representative tools and performed an empirical analysis with the selected tools on Android apps after some adaptation. We evaluated the performance of the tools in terms of their effectiveness and efficiency. Overall, 2,142 unique checked-in secrets were detected, affecting 2,115 Android apps. We further performed a systematic comparison among the tools to analyze the limitations of current checked-in secret detection techniques. Based on our findings, we proposed possible research directions for developing a more effective checked-in secret detection tool specifically for Android apps. The checked-in secrets resulted from our experiments form a benchmark for checked-in secret detection in Android apps. The sanitized datasets are available on our project website (dat, [n. d.]) without leaking any sensitive checked-in secrets.

### v2

@@ 5.2.1 Complimentary Results of the Tools
In contrast, machine-learning-based tools, in general, can adapt to a wider range of cases since they do not require any pre-defined patterns. However, our results suggest that simply providing more training sets cannot yield a significant improvement. The effectiveness of the context model is significantly degraded by the presence of obfuscation. Relying on code snippets as input for the model is proven to be ineffective in our evaluation. This finding suggests the need for a new feature extraction.
5.2.2 Strings as Features for Secret Detection
Table 6: String Group that contains a Google API key
String Group ID Strings
String Group 1 grant_type, refresh_token, client_id, client_secret, refresh_token
String Group 2 account, google, user_id, client_secret, content
String Group 3 builder, google, refresh, token, builder
@@ 5.2.2 Strings as Features for Secret Detection
After analyzing the methods that contain a checked-in secret, we propose to use string groups as features to detect checked-in secrets. A string group refers to all the strings within one single method. In our manual inspection, we observed that string groups that contain checked-in secrets shared some similar patterns. For instance, the names of cloud providers like “Google”or “Firebase ”, along with terms related to secrets such as “user”, “token”, or “key”, frequently appear in these groups. As an example, Table 6 is three string groups that were extracted from different Android apps which contain a Google API key. While the individual strings within these groups are not necessarily identical, they exhibit notable similarity due to the frequent occurrence of certain words (e.g., refresh, token, client_id, etc.).
This suggests that the patterns in string groups could serve as potential indicators of checked-in secrets. Moreover, since string literals are typically not modified by the obfuscation of apps in Google Play Store obfuscation, it can resolve the existing challenge caused by obfuscation for checked-in secret detection in Android apps. We conducted two preliminary studies to test this potential research direction.
@@
Study 1: Statistical Hypothesis Testing To verify the feasibility of this approach, we first conducted a preliminary study with the hypothesis that string groups containing checked-in secrets exhibit similarities to one another and are distinct from those that do not contain any secrets. To test this hypothesis, we conducted a statistical analysis. First, we collected two collections of string groups: Collection Secret included 602 string groups containing Google API keys and secrets obtained by the Three-Layer Filter. we excluded string groups that contain only the checked-in secret without any additional strings. To prevent potential bias in terms of quantity, the second set, Collection NoSecret, included 602 randomly selected string groups that did not contain any checked-in secrets. Methods with only a single string were excluded to ensure meaningful comparison. We then calculated the similarity between string groups using cosine similarity cos, which is better for this task as it works with non-English text, unlike Levenshtein distance lev or Word Movers’ Distance wmd. Cosine similarity measures the cosine of the angle between two character vectors, generating a value between 0 and 1, where 1 indicates identical word vectors and 0 indicates no similarity. The similarity score is calculated using the equation below cos.
$\text{sim}(x,y)=\frac{x\cdot y}{\|x\|\cdot\|y\|}$ (1)
@@
For each string group in Collection Secret, we first transform the string groups into the character vector, in which the number in the vector represents the frequency of a character. Then, we computed the average cosine similarity both within Collection Secret (Comparison-SvS) and between Collection Secret and Collection NoSecret (Comparison-SvN). We need to verify that the two comparisons are statistically different to test our hypothesis. We applied both F-test ftest and Z-test ztest. These tests assessed whether the means and variances of the cosine similarity scores in Comparison-SvS and Comparison-SvN were significantly different. The statistical significance was determined through p-values, with smaller p-values indicating a greater likelihood that the differences in means or variances were not due to chance. Additionally, given the various forms that strings can take in Android apps, we modified Collection Secret and Collection NoSecret into three distinct versions: case-sensitive, case-insensitive, and English Word Extraction. The first two variations considered whether string case impacted the similarity calculations, while the English Word Extraction version isolated all English words using a dictionary englishdict for further analysis.
@@
The result of the two tests is shown in Table 7. The calculated p values of the two tests for the three versions are all very close to 0 and much smaller than the common significance level alpha (0.01) alpha, indicating that differences of Comparison S-S and Comparison S-N are statistically significant. We observed that the difference in the case-sensitive comparison is larger than in other variations. This is likely because ordinary strings tend to use more capital letters. Furthermore, the mean similarity score in Comparison S-S is also higher than in Comparison S-N.
@@
Such test results demonstrate the potential to leverage string groups to detect checked-in secrets in obfuscated Android apps. We plan to further explore this direction to propose a new Machine-Learning-Based checked-in secret detection technique based on string groups.
Table 7: Result of F-test and Z-test
F-Test(p value) Z-Test(p value)
Similarity Score
(Comparison s-s:
Comparison s-n)
@@
English Word $7.6\times 10^{-12}$ 0 0.65 : 0.55
Study 2: Binary Classifiers Using String Groups for Checked-in Secret Detection
To demonstrate the potential of string group analysis, we conducted an additional experiment to actually train machine learning models with string groups.
We first extracted features from the Collection Secret and Collection NoSecret datasets using Term Frequency-Inverse Document Frequency (TF-IDF) and Count Frequency, respectively. These extracted features were then used to train three classification models: Logistic Regression, Naive Bayes, and Support Vector Classification (SVC). Using this experiment, we aim to evaluate the performance of using string groups as features to detect checked-in secrets.
Dataset Pre-processing We used the same dataset as in Study 1, and similar to statistical hypothesis testing, we transformed the two collections into three distinct versions: case-sensitive, case-insensitive, and English word extraction. The case-sensitive and case-insensitive versions examined the impact of string cases. The English word extraction version focused on isolating all English words from the string groups using a dictionary englishdict. Each version was then used to generate collection pairs of the same type, which were combined to construct the complete dataset. As a result, we obtained three distinct datasets, each corresponding to one of the three versions.
Feature Extraction We selected two simple and relatively popular feature extraction techniques used in the domain of text classification: TF-IDF and Count frequency.
TF-IDF is a statistical measure used to evaluate the importance of a word in a document relative to a collection of documents (string groups). It multiplies two following metrics tfidf.
  • Term Frequency (TF): This measures how frequently a term appears in a document. In our case, each string group is treated as a document. It is computed by counting the occurrences of a term within a string group, as shown in Equation 2.
    $tf_{\textit{t,d}}=\text{count(t,d)}$ (2)
  • Inverse Document Frequency (IDF): This assesses the importance of a term within the entire corpus. It is calculated by taking the logarithm of the total number of documents(string groups) divided by the number of documents(string groups) containing the term, as shown in Equation 3.
[…]
Count Frequency simply counts the occurrences of each word in a document(string group) without considering their distribution across the whole dataset.
Both feature extraction methods produce a vector representing the computed value of each term for every string group.
Model Training We first split the datasets into an 80% training set and a 20% testing set. Using the scikit-learn package log; svc; naive, we trained three models (Logistic Regression, Naive Bayes, and SVC) on each dataset separately. For each model, we selected the hyperparameters that achieved the highest training accuracy.
Result We calculated the precision, recall, accuracy and F1 score for each selected model on the different datasets. These metrics were used to evaluate the performance of the models, providing a comprehensive understanding of their effectiveness across the various versions of the datasets. The results of the experiment are listed in Table 8, Table 9, and Table 10, respectively. Each table represents the performance of three models on their corresponding dataset. As shown in the tables, all models demonstrated strong performance, with SVC achieving the highest overall performance, reflected by an average F1-score of 0.97 and an average accuracy of 0.98. Additionally, the two feature extraction techniques evaluated in this study achieved comparable performance, with count frequency performing slightly better than TF-IDF. This is reflected by an average accuracy of 0.96 for both techniques, and average F1-scores of 0.94 for count frequency compared to 0.93 for TF-IDF. However, we observed that the English word extraction version underperformed, which may be due to the fact that some important information, such as URLs and token prefixes, could be ignored during the extraction process. The results highlight that, despite relying on straightforward feature extraction methods and simple models, string group analysis achieves remarkably effective detection performance. This demonstrates the potential of our approach as a lightweight yet powerful solution.
While the results demonstrate the effectiveness of the proposed approach, they are based on a small-scale dataset only composed of Google API keys, which may limit the generalizability and robustness of the findings. This controlled setting was chosen due to the availability of sufficient labeled samples for meaningful training and evaluation. To address this limitation, future work will focus on expanding the dataset to include a broader range of secret types and more diverse application contexts. Furthermore, as discussed previously, the underperformance of the English word extraction version may be attributed to the loss of important information during the pre-processing step. Future work should also focus on developing more advanced pre-processing techniques capable of accurately preserving such information, as well as exploring more complex machine learning models that can better capture these patterns.
Table 8: Result of Three Models on Case Sensitive Dataset
Precision
(TF-IDF
Count Frequency)
Recall
(TF-IDF
Count Frequency)
Accuracy
(TF-IDF
Count Frequency)
F1-Score
(TF-IDF
Count Frequency)
Logistic
Regression
1 0.99 0.99 0.99
1 0.98 0.99 0.99
Naive Bayes 1 0.94 0.99 0.97
1 0.94 0.99 0.97
SVC 1 1 1 1
1 1 1 1
Table 9: Result of Three Models on Case Insensitive Dataset
Precision
(TF-IDF
Count Frequency)
Recall
(TF-IDF
Count Frequency)
Accuracy
(TF-IDF
Count Frequency)
F1-Score
(TF-IDF
Count Frequency)
Logistic
Regression
1 0.73 0.96 0.84
1 0.99 0.99 0.99
Naive Bayes 1 0.94 0.99 0.97
1 0.94 0.99 0.97
SVC 1 1 1 1
1 1 1 1
Table 10: Result of Three Models on English Word Extraction Dataset
Precision
(TF-IDF
Count Frequency)
Recall
(TF-IDF
Count Frequency)
Accuracy
(TF-IDF
Count Frequency)
F1-Score
(TF-IDF
Count Frequency)
Logistic
Regression
1 0.77 0.9 0.87
0.91 0.86 0.9 0.88
Naive Bayes 0.87 0.91 0.9 0.89
0.8 0.92 0.86 0.86
SVC 1 0.85 0.93 0.92
0.97 0.85 0.93 0.91
6 Discussion
@@
Severity of Checked-in Secret Issue in Android apps. Our main study on top-ranked applications demonstrated that checked-in secret leakage is a persistent and non-trivial security concern. To examine whether these findings also apply to long-tail applications, we conducted an additional experiment on a random sample of 386 applications drawn from the 1,565,139 total reported by AppBrain appBrain, achieving a 95% confidence level. The experimental setup was identical to the main study, with LeakScope and the Three-Layer Filter configured to the same set of cloud providers. Manual validation was performed for all detected instances across all tools, following the same procedure as in the original experiment.
In the long-tail sample, LeakScope detected 25 checked-in secrets, consisting of 20 Google API keys, 1 Amazon key, 1 Twitter key, and 2 Facebook keys, all confirmed as true positives. The Three-Layer Filter detected 14 Google API keys, also validated as true positives, but none of the 381 sampled Twitter results were validated as true positives. PassFinder, when applied to the Google API keys detected by the Three-Layer Filter along with their contextual surroundings, did not identify any as valid.
We further analyzed the sampled applications using their download counts and last update dates, as illustrated in Figure 3 and 4. The distribution indicates that a substantial portion of the sample that contains leaked checked-in secrets falls into lower-download categories: 16.67% have fewer than 1,000 installs, 3.33% between 1,000–10,000 installs, and 20.00% between 10,000–100,000 installs, collectively representing over 40% of the sample. Similarly, a majority of the apps were updated recently, showing that secret leakage is not confined to abandoned or outdated apps. Although these results highlight the presence of leakage in long-tail and less frequently downloaded applications, the additional experiment does not necessarily prove that such apps are inherently more dangerous than popular ones. As demonstrated in our original experiment, widely used applications also suffer from secret leakage, sometimes at significant scale. Taken together, these findings suggest that leakage is not restricted to any single segment of the ecosystem: it affects both niche and highly popular apps, including those actively maintained, underscoring that checked-in secret leakage is a pervasive and systemic problem across Android applications.
Figure 3: Number of Apps with Leaked Secrets by Install Range
Figure 4: Number of Apps with Leaked Secrets by Last Updated Time (Monthly)
The same limitations observed in the main study were also present in the long-tail sample. High false negative rates and a strong dependency on pre-defined patterns or API lists remain major challenges. This indicates that the risk of checked-in secret leakage is not limited to top-ranked applications and may be equally prevalent across the broader application ecosystem. Although we could not conduct a larger-scale experiment to precisely measure the relationship between application popularity and security risk due to time constraints, this is an important direction for future research.
Negligence of Checked-in Secret Issue. We reached out to the developers of Android apps affected by the checked-in secret issue. Some developers acknowledged the problem and committed to addressing it, while others acknowledged the issue but asserted that the checked-in secret was safe to expose. While it is true that some checked-in secrets, such as the Firebase API key, are considered safe to expose in code (though not recommended) firebase. Or Twitter Client Secret, which would only become harmful if the associated Twitter Client ID is also leaked, these keys can still pose a security threat when other related information is also exposed. For instance, if project IDs and authentication tokens associated with these checked-in secrets are leaked, they can compromise security. In our experiment, we successfully extracted additional information associated with the Firebase API key that could be harmful when combined, demonstrating the potential risk even for seemingly harmless checked-in secrets.
@@ 7 Threats to Validity
It is also important to acknowledge that our evaluation focused on only three detection tools. However, the selection of these tools was grounded in a rigorous systematic literature review, ensuring that they represent the state-of-the-art approaches within their respective categories. Specifically, each tool was chosen to reflect a different detection methodology, allowing for a comprehensive comparison across diverse detection strategies.
@@ 8 Related work
Checked-in Secret Detection for GitHub. Leakage of checked-in secret issue in the code-sharing platform is a long-studied subject. Besides the tools we selected to run the experiment, there exist other papers that also targeted the checked-in secret issue. Saha et al. reducefalse introduce a machine learning-based detection tool; they designed 20 features for a string to reduce the false positive, and Lounici et al. Lounic examined the performance of various machine learning model when performing checked-in secret detection tasks. Basak et al. assetha developed a framework to detect secret-asset pairs in open-source repository with a focus on database key. There are also some community checked-in secret detection tools. As an example, TruffleHog trufflehog, which was originally designed for GitHub, has also been adapted to the Android apps on the Google Play store. However, the Three-Layer Filter, which utilizes the same methodology, outperforms TruffleHog. Some researchers have also proposed tools that can detect human-generated passwords. Fadi et al.  humanpass utilized neural networks to predict whether a string is a human-generated password. While Lykousas et al. tales examined the developers’ traits when choosing passwords. Those studies all aimed to develop an effective checked-in secret detection tool for open-source projects but did not examine the feasibility of their approaches on Android apps.
@@
Empircal Study on Checked-in Secrets Issue. There have also been several researches that conducted empirical studies on checked-in secret topics. Basak et al. emstudy evaluated the performance of all the available checked-in secret detection tools on a benchmark dataset secretbench that contains 818 Github repositories. They also discussed developers’ challenges about checked-in secret by analyzing 779 questions on Stack Exchange basak2023challenges. Whereas Rahman et al. rahman aimed to investigate the phenomenon of developers disregarding warnings issued by checked-in secret detection tools, often due to the tools’ high false positive rates. Furthermore, Krause et al. krause2022committed analyzed the reason why developers would commit checked-in secrets by interviewing other developers. Naiakshina et al.  Naiakshina_2017 studied how developers deal with checked-in secret storage through an experiment involving 20 developers. However, these empirical analyses have primarily focused on open-source projects, which may not accurately represent the situation in Android apps. We aim to address this limitation in our research.
@@
Security Concerns in Mobile Apps. The popularity of mobile apps has also drawn researchers’ attention. Many studies were conducted to analyze the security issues in mobile apps. PAMDroid pamdroid was developed to investigate the privacy leakage issue caused by the misconfiguration of analytic service. DroidContext DROIDCONTEXT targeted the detection of privacy leakage and is capable of differentiating malicious privacy identification and benign privacy disclosure identification. Bunyakiati et al. secretmanage examined current secret management and handling in the mobile apps development cycle. Whereas Rahat et al. developed OAUTHLINT oauthlint to detect the security mistakes of the Android apps that use OAuth API. our work further investigated the security issue caused by checked-in secret.
9 Conclusion
@@
In this paper, we conducted a literature review to categorize the existing checked-in secret detection techniques, selected three representative tools and performed an empirical analysis with the selected tools on Android apps after some adaptation. We evaluated the performance of the tools based on the number of detected checked-in secrets, false positives and false negatives. Overall, 2,142 unique checked-in secrets were detected, affecting 2,115 Android apps. We further performed a systematic comparison among the tools to analyze the limitations of current checked-in secret detection techniques. Based on our findings, we proposed possible research directions for developing a more effective checked-in secret detection tool specifically for Android apps and conducted two preliminary experiments to support the proposed research directions. The checked-in secrets resulted from our experiments form a benchmark for checked-in secret detection in Android apps. The sanitized datasets are available on our project website data without leaking any sensitive checked-in secrets and the complete dataset can be access upon request.
Disclosures and declarations
Conflict of Interest. The authors declared that they have no conflict of interest.
Data availability. Our code and dataset are available on our project website data. For the dataset, we provided sanitized code snippets and data flow of the detected checked-in secrets. Due to the sensitiveness of the dataset, we did not provide the original Android apps we used in the experiment. The Android apps can be shared to selected researchers and developers upon request.
Funding. This work was supported by Fonds de recherche du Québec(Grant No.2024-NOVA-346499) nova, Natural Sciences and Engineering Research Council of Canada Discovery Grant(Grant No. RGCPIN-2022-03744 and Grant No. DGECR-2022-00378) and Fonds de recherche du Québec-secteur Nature et technologies(Grant No.363482 frqntNewAca).
Ethical Approval. Not Applicable.
Informed Consent. Not Applicable.
Clinical Trial Number. Not Applicable.
Author Contribution. Kevin Li - Conceptualization, Methodology, Data collection, Manual Validation, Formal analysis, Literature Review, Writing. Lin Ling - Data collection, Manual Validation, Literature Review, Writing - review & editing. Jinqiu Yang - Conceptualization, Methodology, Writing review & editing, Supervision. Lili Wei - Conceptualization, Methodology, Writing review & editing, Supervision.